OpenStudy (anonymous):

Help worded question 3 only Please

OpenStudy (anonymous):

@jdoe0001 can you help me please

OpenStudy (jdoe0001):

well, they're really just asking for the inverse function, is all Pedro wants to know when[at what week] he will complete 30 assignments, he just dunnos what the inverse function of f(x) is

OpenStudy (anonymous):

can you solve it for me cause I dont know how?

OpenStudy (jdoe0001):

\(\bf f(x) = y = \cfrac{4x-3}{2}\quad\\ \textit{we get the inverse of it, by }\text{changing the letters}\\ \color{red}{y} = \cfrac{4\color{blue}{x}-3}{2}\quad \textit{inverse is}\quad \color{blue}{x} = \cfrac{4 \color{red}{y}-3}{2}\)

OpenStudy (jdoe0001):

then just solve for "y" on the inverse

OpenStudy (anonymous):

how do you solve for Y

OpenStudy (jdoe0001):

well, there's the assumption you're supposed to know that by now

OpenStudy (jdoe0001):

solve, or "isolate y"

OpenStudy (anonymous):

okay let me try multiple both sides by 2?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

2x=4y-3?

OpenStudy (anonymous):

then what

OpenStudy (jdoe0001):

add 3 to both sides

OpenStudy (anonymous):

2x+3=4y next

OpenStudy (jdoe0001):

divide both sides by 4

OpenStudy (anonymous):

okay 2x+3/4

OpenStudy (jdoe0001):

\(\bf x = \cfrac{4y-3}{2} \implies 2x = 4y-3 \implies 2x+3 = 4y\implies \cfrac{2x+3}{4} = y\)

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

2x+3/4=y

OpenStudy (jdoe0001):

that's what Pedro needs, to plug in the assignments and get the "weeks" it'll take

OpenStudy (anonymous):

do I have to do it for him plug the 30 into the x?

OpenStudy (jdoe0001):

ohhh.. hmm, sure, yes, set x = 30, see how many weeks it takes, or see what "y" is

OpenStudy (anonymous):

okay next one?

OpenStudy (jdoe0001):

\(\bf y= \cfrac{2x+3}{4} \implies y= \cfrac{2(30)+3}{4} \implies y= \cfrac{60+3}{4}\ 15.75 so, 15 weeks and about day and a half

OpenStudy (jdoe0001):

woops

OpenStudy (jdoe0001):

\(\bf y= \cfrac{2x+3}{4} \implies y= \cfrac{2(30)+3}{4} \implies y= \cfrac{60+3}{4}\)

OpenStudy (anonymous):

@jdoe0001 ? You there

OpenStudy (jdoe0001):

she played 14hrs already she's restricted to 4 per week, only for 7 weeks if we make the weeks "x" f(x) = 4x but she already has 14hrs played from last week so f(x) = 4x + 14

OpenStudy (jdoe0001):

well, you may want post anew, so we can all see it and help and revise each other gives more exposure, or more eyes :)